OpenStudy (anonymous):

what is the diagonal of a 30 meters square?

OpenStudy (anonymous):

Same method Use pythagoras theorem

OpenStudy (anonymous):

In this case use p=b as all sides of a square are equal.

OpenStudy (anonymous):

p=b=30meters and find h, which is the diagonal.

OpenStudy (anonymous):

d^2 = 30^2 + 30^2 = 1800 d = sqrt 1800 = 42.43 m

OpenStudy (mertsj):